# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 5.4. Sobolev embedding
We know that the standard Sobolev inequality with $p\in(2,\infty)$
[…]
holds for a function $f$. Now we set
(113) $f=\sqrt{\mathcal{G}_{A_{1}A_{2}A_{3}\cdot\cdot\cdot\cdot A_{N}}\mathcal{G}~_{B_{1}B_{2}B_{3}\cdot\cdot\cdot\cdot B_{N}}\gamma^{A_{1}B_{1}}\gamma^{A_{2}B_{2}}\gamma^{A_{3}B_{3}}\cdot\cdot\cdot\cdot\gamma^{A_{N}B_{N}}+\delta}$
with $\delta>0$ to yield
[…]
since $f$ is a function and the covariant derivative is compatible with the metric $\gamma$. Substituting this inequality in (112) and taking the limit $\delta\to 0$ yields the desired gauge-invariant Sobolev inequality. ∎
Proposition 5.9.
Let $(S,\gamma)$ be a Riemannian 22-manifold with the isoperimetric constant $I(S)$. Then the following Sobolev embedding holds for any $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S)$
[…]
for any $p\in(2,\infty)$.
@@
A calculation similar to the previous one and the standard $L^{\infty}-L^{p}$ Sobolev inequality on the Riemannian manifold $(S,\gamma)$ yield the result. ∎
@@
The two inequalities above, together with Propositions 5.1-5.4, allow us to control the $L^{2}$-norm of $\mathcal{G}$ in terms of its $H^{2}$-norm. Following the area estimates, we have $Area(S_{u,\underline{u}})\approx u^{2}$. Therefore, we obtain the following important inequality.
Proposition 5.10.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), the following Sobolev embedding holds for any $\mathcal{G}\in\Gamma(^{N}\otimes T^{*}S_{u,\underline{u}}):$
[…]
which, in scale-invariant norms, reads
[…]
Proof.
Substitute $p=4$ in the previous proposition (114) and estimate the right hand side by means of (111), the estimate on the area given by Proposition 5.4 and on the isoperimetric constant of $S_{u,\underline{u}}$ given by Proposition 110. ∎
@@ 5.5. Estimates on the Ricci coefficients
Proposition 5.11.
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\left(1+\hskip 1.42262pt\underline{\mathcal{R}}[\rho]\right).$
@@
Using the commutation formula 2 and the notation of Section 4.6, we have, for a general $i$:
(117) $\begin{split}\nabla_{3}\nabla^{i}\omega+\frac{i}{2}\mbox{tr}\underline{\chi}\nabla^{i}\omega=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega\end{split}$
Note that since $\omega$ is not a section of the vector bundle $^{k}\otimes T^{*}S\otimes P_{Ad,\mathfrak{g}}$, the last term on the right-hand side of 2 does not appear. Passing to scale-invariant norms, we get
@@
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
@@
For $0\leq i\leq N+4$, the first term, by virtue of the fact that $\Omega\equiv 1$ initially, vanishes. The second term can be bounded, using Hölder’s inequality, by $\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\underline{\mathcal{R}}[\rho].$ The third, fourth and fifth terms can be bounded above by
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\leq\frac{a\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert^{2}}\leq\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}.$
@@
(118) $\begin{split}&\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{4}}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{3}{4}}\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert}\lesssim\frac{a^{\frac{1}{4}}}{\lvert u\rvert^{\frac{1}{2}}}.\end{split}$
@@
(119) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert^{2}}{a}\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{\Gamma^{3}}{\lvert u\rvert}\lesssim\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\lesssim 1.$
@@
(120) $\lesssim$ $\int_{u_{\infty}}^{u}\int_{u_{\infty}}^{u}\frac{a^{2}}{|u^{{}^{\prime}}|^{2}}\frac{1}{|u^{{}^{\prime}}|^{2}}\frac{|u^{{}^{\prime}}|}{a^{\frac{1}{4}}}\Gamma^{3}du^{{}^{\prime}}\lesssim\frac{a^{\frac{7}{4}}\Gamma^{3}}{|u|^{2}}\lesssim 1$
@@
$\sum_{i\leq N+4}\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1,\hskip 5.69054pt\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\mathcal{R}[\alpha]+1.$
@@
The proof is exactly similar to that of [11]. ∎
@@
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\leq\mathcal{R}[\rho]+1.$
@@
As before, the schematic product of the Yang-Mills components is gauge-invariant. Using the commutation formula 4.1, we have, for a general $i$:
(121) $\begin{split}\nabla_{4}\nabla^{i}\underline{\omega}=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\psi_{g}\\ &++\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\end{split}$

### v4

@@ 4.3. Sobolev embedding
We use the scalar Sobolev inequality on $(S,\gamma)$ in the formulation controlled by the isoperimetric constant: for each $p\in(2,\infty)$ there exists $C_{p}$ such that for every $f\in W^{1,2}(S)$,
(137) $\text{Area}(S)^{-\frac{1}{p}}\,\|f\|_{L^{p}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\Big(\|\nabla f\|_{L^{2}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|f\|_{L^{2}(S)}\Big).$
(See, e.g., the Federer–Fleming/Maz’ya isoperimetric–Sobolev inequality specialized to dimension $2$.)
To deduce (136) from (137), we apply (137) to the regularized norm of $\mathcal{G}$. For $\delta>0$ define the Lipschitz function
$f_{\delta}:=\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}\,,\qquad|\mathcal{G}|_{\gamma}^{2}:=\gamma^{A_{1}B_{1}}\cdots\gamma^{A_{N}B_{N}}\,\mathcal{G}_{A_{1}\cdots A_{N}}\,\mathcal{G}_{B_{1}\cdots B_{N}}.$
Since $f_{\delta}\in W^{1,2}(S)$ and $S$ is compact, (137) applies to $f_{\delta}$.
Estimate of $\nabla f_{\delta}$. By metric compatibility $\nabla\gamma=0$ and the Leibniz rule,
$\nabla(|\mathcal{G}|_{\gamma}^{2})=2\langle\mathcal{G},\nabla\mathcal{G}\rangle_{\gamma},$
where $\langle\cdot,\cdot\rangle_{\gamma}$ denotes the pointwise inner product induced by $\gamma$ on the relevant tensor bundle. Hence, in the weak sense (and pointwise a.e. since $f_{\delta}$ is Lipschitz),
$\nabla f_{\delta}=\frac{1}{2}(|\mathcal{G}|_{\gamma}^{2}+\delta)^{-1/2}\,\nabla(|\mathcal{G}|_{\gamma}^{2})=\frac{\langle\mathcal{G},\nabla\mathcal{G}\rangle_{\gamma}}{\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}}.$
By Cauchy–Schwarz,
$|\nabla f_{\delta}|\leq\frac{|\mathcal{G}|_{\gamma}\,|\nabla\mathcal{G}|_{\gamma}}{\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}}\leq|\nabla\mathcal{G}|_{\gamma},$
and therefore
(138) $\|\nabla f_{\delta}\|_{L^{2}(S)}\leq\|\nabla\mathcal{G}\|_{L^{2}(S)}.$
Moreover, since $f_{\delta}\geq|\mathcal{G}|_{\gamma}$ and $f_{\delta}\downarrow|\mathcal{G}|_{\gamma}$ pointwise as $\delta\to 0$,
(139) $\|f_{\delta}\|_{L^{2}(S)}\to\|\mathcal{G}\|_{L^{2}(S)},\qquad\|f_{\delta}\|_{L^{p}(S)}\to\|\mathcal{G}\|_{L^{p}(S)}\quad\text{for every }p\in[1,\infty),$
by monotone convergence (or dominated convergence on the compact manifold $S$).
Applying (137) to $f_{\delta}$ and using (138) gives
$\text{Area}(S)^{-\frac{1}{p}}\,\|f_{\delta}\|_{L^{p}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\Big(\|\nabla\mathcal{G}\|_{L^{2}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|f_{\delta}\|_{L^{2}(S)}\Big).$
Letting $\delta\to 0$ and using (139) yields (136). ∎
Proposition 4.8 ($L^{\infty}$ Sobolev embedding in terms of the isoperimetric constant).
Let $(S,\gamma)$ be a smooth compact Riemannian 22–manifold, with isoperimetric constant $I(S)$ and area $\text{Area}(S)$. Fix $p\in(2,\infty)$. Then there exists a constant $C_{p}<\infty$, depending only on $p$, such that the following holds.
For every smooth tensorfield $\mathcal{G}$ of type $(0,N)$ on $S$ (in particular $\mathcal{G}\in\Gamma(\otimes^{N}T^{*}S)$, or more generally any tensor bundle endowed with the norm induced by $\gamma$),
(140) $\|\mathcal{G}\|_{L^{\infty}(S)}\;\leq\;C_{p}\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\Big(\|\nabla\mathcal{G}\|_{L^{p}(S)}+\text{Area}(S)^{-\frac{1}{2}}\,\|\mathcal{G}\|_{L^{p}(S)}\Big).$
@@
We reduce to the scalar inequality and then pass to tensorfields by a gauge–invariant regularization argument, as in Proposition 4.7.
First control scalar $W^{1,p}\hookrightarrow L^{\infty}$ with isoperimetric control in 22-dimensions. Let $f\in W^{1,p}(S)$ with $p>2$. By the isoperimetric inequality, the $L^{1}$–Sobolev inequality holds with constant controlled by $\max\{1,I(S)\}^{1/2}$:
(141) $\|f-\bar{f}\|_{L^{2}(S)}\;\leq\;C\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\|\nabla f\|_{L^{1}(S)},\qquad\bar{f}:=\text{Area}(S)^{-1}\int_{S}f.$
Interpolating between $L^{1}$ and $L^{p}$ norms of $\nabla f$ and using Hölder’s inequality yields, with a constant depending only on $p$,
(142) $\|f-\bar{f}\|_{L^{2}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\,\|\nabla f\|_{L^{p}(S)}.$
Next, using the standard Morrey embedding in dimension $2$ on compact manifolds (which can be proved by covering $S$ with finitely many coordinate charts and reducing to the Euclidean Morrey inequality), we have
(143) $\|f-\bar{f}\|_{L^{\infty}(S)}\;\leq\;C_{p}\Big(\|\nabla f\|_{L^{p}(S)}+\text{Area}(S)^{-1/2}\|f-\bar{f}\|_{L^{2}(S)}\Big).$
Combining (142) and (143) gives
(144) $\|f-\bar{f}\|_{L^{\infty}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\,\|\nabla f\|_{L^{p}(S)}.$
Finally, control $\bar{f}$ by Hölder:
$|\bar{f}|\leq\text{Area}(S)^{-1}\|f\|_{L^{1}(S)}\leq\text{Area}(S)^{-1/p}\|f\|_{L^{p}(S)}.$
Thus, from (144),
(145) $\|f\|_{L^{\infty}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\Big(\|\nabla f\|_{L^{p}(S)}+\text{Area}(S)^{-\frac{1}{2}}\|f\|_{L^{p}(S)}\Big).$
Now let $\mathcal{G}$ be a smooth tensorfield. For $\delta>0$ define
$f_{\delta}:=\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}\in W^{1,p}(S).$
As in the proof of Proposition 4.7, metric compatibility and the chain rule yield (a.e.)
$\nabla f_{\delta}=\frac{\langle\mathcal{G},\nabla\mathcal{G}\rangle_{\gamma}}{\sqrt{|\mathcal{G}|_{\gamma}^{2}+\delta}},\qquad|\nabla f_{\delta}|\leq|\nabla\mathcal{G}|_{\gamma},$
hence
(146) $\|\nabla f_{\delta}\|_{L^{p}(S)}\leq\|\nabla\mathcal{G}\|_{L^{p}(S)}.$
Moreover, $f_{\delta}\downarrow|\mathcal{G}|_{\gamma}$ pointwise and monotonically as $\delta\to 0$, so
(147) $\|f_{\delta}\|_{L^{p}(S)}\to\|\mathcal{G}\|_{L^{p}(S)},\qquad\|f_{\delta}\|_{L^{\infty}(S)}\to\|\mathcal{G}\|_{L^{\infty}(S)}.$
Apply the scalar inequality (145) to $f_{\delta}$ and use (146):
$\|f_{\delta}\|_{L^{\infty}(S)}\;\leq\;C_{p}\,\big(\max\{1,I(S)\}\big)^{\frac{1}{2}}\,\text{Area}(S)^{\frac{1}{2}-\frac{1}{p}}\Big(\|\nabla\mathcal{G}\|_{L^{p}(S)}+\text{Area}(S)^{-\frac{1}{2}}\|f_{\delta}\|_{L^{p}(S)}\Big).$
Letting $\delta\to 0$ and using (147) yields (140). ∎
@@
The two inequalities above, together with Propositions 4.1-LABEL:34, allow us to control the $L^{2}$-norm of $\mathcal{G}$ in terms of its $H^{2}$-norm. Following the area estimates, we have $Area(S_{u,\underline{u}})\approx u^{2}$. Therefore, we obtain the following important inequality.
Proposition 4.9 (Sobolev embedding on $S_{u,\underline{u}}$).
Assume the hypotheses of Theorem 2.2 and the bootstrap assumptions (118). Let
$S_{u,\underline{u}},\qquad u\in[u_{\infty},-a],\ \underline{u}\in[0,1],$
be the 22–spheres of the canonical double–null foliation with induced metric $\gamma=\gamma(u,\underline{u})$, Levi–Civita connection $\nabla$, and area form $d\mu_{\gamma}$. Then for every smooth tensorfield $\mathcal{G}\in\Gamma(\otimes^{N}T^{*}S_{u,\underline{u}})$ one has the uniform Sobolev estimate
(148) $\|\mathcal{G}\|_{L^{\infty}(S_{u,\underline{u}})}\;\lesssim\;\sum_{I=0}^{2}\big\||u|^{I-1}\nabla^{I}\mathcal{G}\big\|_{L^{2}(S_{u,\underline{u}})},$
where the implicit constant is universal (independent of $u,\underline{u},a$) and depends only on the fixed bootstrap constants.
Equivalently, in the scale–invariant norms associated to the spherical scale $|u|$ (and the parameter $a$ fixed in Theorem 2.2)),
(149) $\|\mathcal{G}\|_{L^{\infty}_{sc}(S_{u,\underline{u}})}\;\lesssim\;\sum_{I=0}^{2}\big\|(a^{1/2}\nabla)^{I}\mathcal{G}\big\|_{L^{2}_{sc}(S_{u,\underline{u}})}.$
Proof sketch.
Fix $(u,\underline{u})$. Apply the $L^{\infty}$ Sobolev inequality of Proposition 4.8 to $(S_{u,\underline{u}},\gamma)$ with $p=4$:
$\|\mathcal{G}\|_{L^{\infty}(S_{u,\underline{u}})}\;\lesssim\;\big(\max\{1,I(S_{u,\underline{u}})\}\big)^{1/2}\,\text{Area}(S_{u,\underline{u}})^{1/4}\Big(\|\nabla\mathcal{G}\|_{L^{4}(S_{u,\underline{u}})}+\text{Area}(S_{u,\underline{u}})^{-1/2}\|\mathcal{G}\|_{L^{4}(S_{u,\underline{u}})}\Big).$
Next, estimate the $L^{4}$–terms by the $L^{2}$–Sobolev inequality of Proposition 4.7 with $p=4$, yielding control by $\|\nabla^{2}\mathcal{G}\|_{L^{2}}+\text{Area}(S_{u,\underline{u}})^{-1/2}\|\nabla\mathcal{G}\|_{L^{2}}+\text{Area}(S_{u,\underline{u}})^{-1}\|\mathcal{G}\|_{L^{2}}$. Finally, use the geometric bounds provided by Proposition 4.3 (area comparability $\text{Area}(S_{u,\underline{u}})\sim|u|^{2}$) and Proposition 4.6 (uniform isoperimetric control $I(S_{u,\underline{u}})\lesssim 1$) to rewrite the resulting estimate in the form (148). The scale-invariant formulation (149) is the same inequality expressed in the rescaled norms by factoring out the natural spherical scale $|u|$ (equivalently, the $a^{1/2}\nabla$ normalization used throughout the bootstrap). ∎
@@ 4.4. Estimates on the Ricci coefficients
Proposition 4.10 (Scale–invariant $L^{2}$ control of $\nabla\omega$).
Assume the hypotheses of Theorem 2.2 and the bootstrap assumptions (118). Then, for all
$u\in[u_{\infty},-a],\qquad\underline{u}\in[0,\epsilon],$
the lapse/vorticity coefficient $\omega$ satisfies the scale–invariant estimate
(150) $\sum_{i\leq N+4}\,\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\;\lesssim\;\frac{a^{1/2}}{|u|}\;+\;\frac{a^{1/2}}{|u|}\,\underline{\mathcal{R}}[\rho].$
The implicit constant depends only on the fixed bootstrap constants and on $N$.
@@
Using the commutation formula 8 and the notation of Section 3.6, we have, for a general $i$:
(151) $\begin{split}\nabla_{3}\nabla^{i}\omega+\frac{i}{2}\mbox{tr}\underline{\chi}\nabla^{i}\omega=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\mbox{tr}\underline{\chi}\nabla^{i_{4}}\omega\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\underline{\hat{\chi}},\mbox{tr}\underline{\chi})\nabla^{i_{4}}\omega\end{split}$
Note that since $\omega$ is not a section of the vector bundle $^{k}\otimes T^{*}S\otimes P_{Ad,\mathfrak{g}}$, the last term on the right-hand side of 8 does not appear. Passing to scale-invariant norms, we get
@@
$+\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}$
@@
For $0\leq i\leq N+4$, the first term, by virtue of the fact that $\Omega\equiv 1$ initially, vanishes. The second term can be bounded, using Hölder’s inequality, by $\frac{a^{\frac{1}{2}}}{\lvert u\rvert^{\frac{1}{2}}}\underline{\mathcal{R}}[\rho].$ The third, fourth, and fifth terms can be bounded above by
$\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\leq\frac{a^{\frac{3}{2}}\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert^{3}}.$
@@
(152) $\begin{split}&\int_{u_{\infty}}^{u}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{(a^{\frac{1}{2}})^{i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\underline{\hat{\chi}},\widetilde{\mbox{tr}\underline{\chi}})\nabla^{i_{4}}\omega}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert}{a^{\frac{1}{2}}}\frac{\Gamma^{2}}{\lvert u^{\prime}\rvert}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\hskip 1.42262pt\Gamma^{2}}{\lvert u\rvert^{3}}.\end{split}$
@@
(153) $\lesssim$ $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\frac{\lvert u^{\prime}\rvert^{2}}{a}\frac{\Gamma^{3}}{\lvert u^{\prime}\rvert^{2}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma^{3}}{\lvert u\rvert^{2}}.$
@@
(154) $\lesssim$ $\int_{u_{\infty}}^{u}\int_{u_{\infty}}^{u}\frac{a^{2}}{|u^{{}^{\prime}}|^{2}}\frac{1}{|u^{{}^{\prime}}|^{2}}\frac{|u^{{}^{\prime}}|}{a^{\frac{1}{2}}}\frac{a^{\frac{1}{2}}}{|u^{{}^{\prime}}|}\Gamma^{3}du^{{}^{\prime}}\lesssim\frac{a^{2}\Gamma^{3}}{|u|^{3}}.$
Therefore collecting every terms and by boot-strap
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\omega}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{\lvert u\rvert}+\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\underline{\mathcal{R}}[\rho].$
This completes the proof. ∎
@@
$\sum_{i\leq N+4}\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1,\hskip 5.69054pt\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\hat{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim\frac{1}{a^{\frac{1}{2}}}(\mathcal{R}[\alpha]+1).$
@@
The proof is exactly similar to that of [12]. The point to note here is that $\hat{\chi}$ and $\alpha$ have same scaling. ∎
@@
$\sum_{i\leq N+4}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\omega}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\leq\frac{1}{a^{\frac{1}{2}}}\mathcal{R}[\rho]+\frac{1}{a^{\frac{1}{2}}}.$
@@
As before, the schematic product of the Yang-Mills components is gauge-invariant. Using the commutation formula 3.1, we have, for a general $i$:
(155) $\begin{split}\nabla_{4}\nabla^{i}\underline{\omega}=&\nabla^{i}\rho+\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}\rho+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}\psi_{g}\nabla^{i_{4}}(\hat{\chi},\omega)\\ &++\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\\ &+\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\underline{\omega}\end{split}$